# CharSS: Character-Level Transformer Model for Sanskrit Word Segmentation

### v1

@@ Abstract
Subword tokens in Indian languages inherently carry meaning, and isolating them can enhance NLP tasks, making sub-word segmentation a crucial process. Segmenting Sanskrit and other Indian languages into subtokens is not straightforward, as it may include sandhi, which may lead to changes in the word boundaries. We propose a new approach of utilizing a (Character-level Transformer model for Sanskrit Word Segmentation (CharSS). We perform experiments on three benchmark datasets to compare the performance of our method against existing methods. On the UoH+SandhiKosh dataset, our method outperforms the current state-of-the-art system by an absolute gain of 6.72 points in split prediction accuracy. On the hackathon dataset our method achieves a gain of 2.27 points over the current SOTA system in terms of perfect match metric. We also propose a use-case of Sanskrit-based segments for a linguistically informed translation of technical terms to lexically similar low-resource Indian languages. In two separate experimental settings for this task, we achieve an average improvement of 8.46 and 6.79 chrF++ scores, respectively.
CharSS: Character-Level Transformer Model for Sanskrit Word Segmentation
Krishnakant Bhattthanks: Equal Contribution, Karthika N J11footnotemark: 1, Ganesh Ramakrishnan, and Preethi Jyothi Indian Institute of Technology Bombay, India {kkbhatt, karthika, ganesh, pjyothi}@cse.iitb.ac.in
@@ 1 Introduction
Compound words are formed by combining two or more meaningful subwords. In Indian languages, compounds may be formed either through simple concatenation without boundary changes or by following sandhi rules, resulting in boundary modifications. The process of decompounding a compound Sanskrit word involves segmenting it into smaller, meaningful lexical units. Existing methods used for the Sanskrit Word Segmentation (SWS) [^1] We use the term segmentation for the task of splitting a compound word into its meaningful constituents. task can be roughly classified into two categories: tackling the broader task of SWS and sandhi splitting-specific techniques. The former includes works like Sanskrit Heritage Reader (SHR) Gérard (2003); Sriram et al. (2023), which is a lexicon-driven shallow parser. Hellwig and Nehrdich (2018a) processes compound sandhi words at the character level using recurrent and convolutional neural networks. Sandhan et al. (2022) presents TransLIST, integrating a module that appends additional latent information from SHR to the input sequence. It also employs a soft masked attention mechanism to prioritize relevant subword candidates and incorporates a path ranking algorithm to mitigate erroneous predictions. Alternately, Aralikatte et al. (2018) proposes a dual-decoder approach where the first decoder identifies the location for the sandhi split (sandhivicchēda) [^2] We follow ISO-15919 script to mention Roman translations of Indian language text for better readability., and the second decoder predicts the segmented output. Similarly, Dave et al. (2021) applies an RNN encoder-decoder-based two-stage methodology to predict the location and final splits.
In this work, we explore the efficacy of byte/character-level Transformer models for the task of sandhi splitting. To the best of our knowledge, we are the first to explore an entirely character-level encoder-decoder model based on the Transformer architecture for this task. Byte- and character-level models are known for their robustness in tasks sensitive to variations in spelling and pronunciation. We believe this modeling framework to be particularly well-suited to a task such as sandhi splitting.
@@
We also propose a use case of Sanskrit-based segmented morphemes for a linguistically informed translation of technical dictionary terms. Kunchukuttan and Bhattacharyya (2016) shows the importance of subword segmentation and lexical similarity of languages in the translation task. In this paper, we introduce a use case of Sanskrit-based sub-word level segmentation in word and phrase-level translation of academic/technical terminologies to leverage the large overlap of vocabulary among Indian languages.
Our main contributions are:
@@
  • Through comprehensive experiments we show the efficacy of our proposed methodologies. We test our methodology on three benchmark datasets viz., UoH+SandhiKosh, SIGHUM, and Hackthon Datasets for SWS. Similarly, we test and compare our technical term translation method for multiple low-resource languages (Section 3).
@@ 2.2 Technical Term Translation
Motivation to use Hindi data to generate Sanskrit-based segments: There is a significant under-representation of digital resources for all other Indian languages compared to Hindi. Appendix A shows details of this digital data divide. English-Hindi human-translated data is readily available for the domains we considered in this work. We obtained Sanskrit-based sub-word tokens from the available Hindi data for over 76% of training and test instances. Furthermore, a word in one language may have several different translations in another language, depending on the context of usage. Providing the augmented input helps disambiguate the domain of the word. We provide a detailed analysis to support this argument in Appendix A.4.
@@ 3 Experiments and Results
3.1 Data and Metrics
For the SWS task, following Dave et al. (2021) and Sandhan et al. (2022), we use three publicly available benchmark datasets, UoH corpus [^3] https://sanskrit.uohyd.ac.in/Corpus/ combined with the SandhiKosh dataset Bhardwaj et al. (2018), SIGHUM dataset Krishna et al. (2017), and hackathon dataset Krishnan et al. (2020). These datasets are carefully curated subsets of a larger corpus DCS Hellwig (2010). The UoH corpus+SandhiKosh dataset has 62273 and 15569 instances as train and test sets. For this dataset, we apply the pruning technique mentioned in Dave et al. (2021) to filter out invalid instances. The size of the training, validation, and test sets for the SIGHUM dataset are 97000, 3000, and 4200, respectively, and for the hackathon dataset, it is 90000, 10332, and 9963, respectively. Contemporary deep-learning methodologies have demonstrated enhanced performance when utilizing the SLP1 script for Sanskrit. Consequently, we have prepared all datasets in the SLP1 script to leverage these performance improvements. We use word-level accuracy as the evaluation metric for the SWS task. To compare against Sandhan et al. (2022), we also calculate sentence level perfect match (PM) for SIGHUM and hackathon datasets.
Model LPA SPA
JNU - 8.1
UoH - 47.2
INRIA - 59.9
$D$-RNN 95.0 79.5
Sandhi Prakarana 92.3 86.8
ByT5 97.2 93.5
Model SIGHUM Hackathon
P R F PM P R F PM
rcNN-$S$ 96.86 96.83 96.84 87.08 96.40 95.15 95.77 77.62
TransLIST 98.80 98.93 98.86 93.97 97.78 97.44 97.61 85.47
TransLIST - - - 86.10 - - - -
ByT5 98.68 98.42 98.53 93.78 97.58 97.71 97.63 87.7
For the technical term translation task, we utilize the technical bilingual dictionary datasets provided by NJ et al. (2024) which is a dataset curated from CSTT  [^4] https://cstt.education.gov.in/en dictionaries. The dataset consists of word-level translations from English to 6 Indian languages across 3 domains, viz., administrative, biotechnology, and chemistry, and has 9094 terms in the training data and 1285 in the test data for all domains combined. We obtained Sanskrit-based inputs for all data instances by applying our approach of generating Sanskrit-based additional inputs. We use chrF++ Popović (2017) as the evaluation metric for all the experiments under this task.
@@ 3.2 Experiments on the SWS Task
We utilize the pre-trained checkpoint of the base variant of the ByT5 model available via Huggingface  [^5] https://huggingface.co/google/byt5-base and fine-tune it over the UoH+SandhiKosh, SIGHUM dataset, and hackathon datasets as three separate experiments.
@@
Baselines.
For the experiments performed over the UoH+SandhiKosh dataset, we compare our method against Sandhi Prakarana Dave et al. (2021), DD-RNN Aralikatte et al. (2018), and 3 sandhi spitter tools viz (i) JNU Splitter Sachin (2007), (ii) UoH Splitter Kumar et al. (2010), and (iii) INRIA Sanskrit Heritage Reader Huet (2003); Goyal and Huet (2013). We reproduce and report the scores reported by Dave et al. (2021). For DD-RNN and the 3 sandhi tools, we report the scores reported in Aralikatte et al. (2018) and Dave et al. (2021). For the experiments performed over the SIGHUM and hackathon datasets, we compare our method against TransLIST Sandhan et al. (2022) and rcNN-SS Hellwig and Nehrdich (2018b).
@@ Results.
Tables 1 and 2 report the performance of our methodology compared with the baselines over the respective datasets. Table 1 shows that our methodology outperforms all other baselines in terms of both Location Prediction Accuracy (LPA) and Split Prediction Accuracy (SPA) with absolute gains of 4.86 and 6.72, respectively, on the UoH+SandhiKosh dataset. TransLIST Sandhan et al. (2022) utilizes a set of potential split candidates from SHR (referred to as LIST in their paper), which provides additional linguistic information for segmentation. Our model is not linguistically informed like this as we feed only the compound word to the model. Hence, our method is not strictly comparable with the results shown in row 2 of Table 2. Nevertheless, our method outperforms all other models on three out of four evaluation metrics when tested on hackathon dataset. On SIGHUM dataset, our method achieves competitive scores. Sandhan et al. (2022) also reported the performance of their model without the LIST module, as shown in row 3 (TransLIST). The model without the LIST step is more comparable to our setting and we outperform this result as well, while failing to outperform the scores in row 2. As a separate experiment, we provide SHR input to our model for SIGHUM data which outperforms TransLIST on PM metric achieving a PM score of 94.31.
3.3 Experiments on the Technical Term Translation Task
For this task, we have two experimental settings, both formulated as text-to-text translation. In the first setting, we train and test the NMT model NLLB Costa-jussà et al. (2022) over all 6 language pairs across 3 domains. As a separate study, in the second setting, we train the model on Hindi, Gujarati, and Tamil across 3 domains, and test it over the Marathi, Kannada, and Odia across the same domains. This setting is considered to test the model’s performance in a zero-shot setting. In the baseline configuration for this task, the model is fed with English input only. In the configuration corresponding to the proposed method, the English input is augmented with additional Sanskrit-based input prepared as discussed in 2.2. We utilize the pre-trained 1.3B parameter checkpoint of the NLLB model available via Huggingface [^6] https://huggingface.co/facebook/nllb-200-1.3B and fine-tune it over the technical domain dictionary data for both experimental settings.
Results.
Table 3 reports the comparison of chrF++ scores obtained by finetuning the NMT model with English-only input (NLLB) and with augmented input (NLLB+Sanskrit) under the first experimental setting. In Appendix A.3 we analyze the performance of the model with and without additional input in a zero-shot setting. Across experiments, there’s a consistent performance gain with the lexically informed input. Our method archives an average improvement of 8.46 chrF++ scores. We also provide a detailed post-hoc analysis of the predictions in Appendix A.4
4 Conclusion
Limitations
Appendix A Appendix
A.1 Normalisation
A.2 sandhi
A.3 Zero-Shot Translation
@@ A.4 Post hoc analysis
We make a detailed analysis of technical terms’ translation results by a comparative study of the outputs in both the input settings, i.e., with and without the Sanskrit-based augmented output.
@@
[…]
Dataset Hindi Gujarati Kannada Tamil Marathi Odiya
IndicCorp v2 (#tokens in Millions) 6107 901 875 476 795 122
CulturaX Nguyen et al. (2024) (in %) 0.27 <0.05 <0.05 0.07 <0.05 <0.05
Table 6: Disparity in Language Coverage in different datasets

### v2

# LEVOS: Leveraging Vocabulary Overlap with Sanskrit to Generate Technical Lexicons in Indian Languages

@@ Abstract
Translating technical terms into lexically similar, low-resource Indian languages remains a challenge due to limited parallel data and the complexity of linguistic structures. We propose a novel use-case of Sanskrit-based segments for linguistically informed translation of such terms, leveraging subword-level similarity and morphological alignment across related languages. Our approach uses character-level segmentation to identify meaningful subword units, facilitating more accurate and context-aware translation. To enable this, we utilize a Character-level Transformer model for Sanskrit Word Segmentation (CharSS), which addresses the complexities of sandhi and morpho-phonemic changes during segmentation. We observe consistent improvements in two experimental settings for technical term translation using Sanskrit-derived segments, averaging 8.46 and 6.79 chrF++ scores, respectively. Further, we conduct a post hoc human evaluation to verify the quality assessment of the translated technical terms using automated metrics. This work has important implications for the education field, especially in creating accessible, high-quality learning materials in Indian languages. By supporting the accurate and linguistically rooted translation of technical content, our approach facilitates inclusivity and aids in bridging the resource gap for learners in low-resource language communities.
LEVOS: Leveraging Vocabulary Overlap with Sanskrit to Generate Technical Lexicons in Indian Languages
Karthika N J, Krishnakant Bhatt, Ganesh Ramakrishnan, and Preethi Jyothi Indian Institute of Technology Bombay, India {karthika, kkbhatt, ganesh, pjyothi}@cse.iitb.ac.in
@@ 1 Introduction
English is the most widely used language in academic books and as a medium of instruction worldwide. India has 22 official languages in addition to English. Over the years, works like (Hudelson, 1987) have established the power of language in students’ learning. Aligned with these studies, the Indian Government has proposed several changes to the education system, among which multilingual knowledge dissemination assumes an important role. The proposal involves introducing regional languages at various levels of education. Carrying out this proposal leads to massive resource requirements like textbook translations and content creation. With the limited technical content availability in non-English languages, especially at the higher education level, translating technical terms from English to other languages is a challenging task that needs to be addressed.
@@
Maheshwari et al. (2024) presents the importance of domain-specific lexicon generation, especially catering to the technical domains, and its importance for translation tasks with low-resource languages as the target. Kunchukuttan and Bhattacharyya (2016) shows the importance of subword segmentation and lexical similarity of languages in the translation task. Additionally, Sanskrit language is known to be a lexically rich, flexible, and well-structured language with the potential to create meaningful new words easily. In this paper, we introduce a use case of Sanskrit-based sub-word level segmentation in word and phrase-level translation of academic/technical terminologies to leverage the large overlap of vocabulary among Indian languages. For the generation of technical terms in low-resource regional languages, we propose to utilize the high vocabulary overlap of Indo-Aryan and Dravidian languages with Sanskrit, thereby performing a lexically informed translation.
Compound words are formed by combining two or more meaningful subwords. In Indian languages, compounds may be formed either through simple concatenation without boundary changes or by following sandhi rules, resulting in boundary modifications. Decompounding a compound Sanskrit word involves segmenting it into smaller, meaningful lexical units. Existing methods used for the Sanskrit Word Segmentation (SWS) [^1] We use the term segmentation for the task of splitting a compound word into its meaningful constituents. task can be roughly classified into two categories: tackling the broader task of SWS and sandhi splitting-specific techniques. The former includes works like Gérard (2003); Sriram et al. (2023), a lexicon-driven shallow parser. Hellwig and Nehrdich (2018) processes compound sandhi words at the character level using recurrent and convolutional neural networks. Sandhan et al. (2022) presents TransLIST, integrating a module that appends additional latent information from SHR to the input sequence. It also employs a soft masked attention mechanism to prioritize relevant subword candidates and incorporates a path ranking algorithm to mitigate erroneous predictions. Alternately, Aralikatte et al. (2018) proposes a dual-decoder approach where the first decoder identifies the location for the sandhi split (sandhivicchēda) [^2] We follow ISO-15919 script to mention Roman translations of Indian language text for better readability., and the second decoder predicts the segmented output. Similarly, Dave et al. (2021) applies an RNN encoder-decoder-based two-stage methodology to predict the location and final splits. Nehrdich et al. (2024) presented a new language model pre-trained for Sanskrit and further fine-tuned and utilised the model for various downstream tasks including word segmentation, lemmatization and morphosyntactic tagging tasks. We use a similar architecture in our CharSS model, to generate the word-splits.
Our main contributions through the paper are:
@@
  • Through comprehensive experiments, we show the efficacy of our proposed methodologies. We test CharSS on three benchmark datasets for SWS. Similarly, we experiment with our technical term translation process for multiple low-resource languages, generating better-quality technical lexicons in the target languages (Section 3).
@@ 2.2 Technical Term Translation
Motivation to use Hindi data to generate Sanskrit-based segments:
There is a significant under-representation of digital resources for all other Indian languages compared to Hindi. Appendix A shows details of this digital data divide. English-Hindi human-translated data is readily available for the domains we considered in this work. We obtained Sanskrit-based sub-word tokens from the available Hindi data for over 76% of training and test instances. Furthermore, a word in one language may have several different translations in another language, depending on the context of usage. Providing the augmented input helps disambiguate the domain of the word. See Section 3.3 for a detailed analysis supporting this argument.
@@ 3 Experiments and Results
For the technical term translation task, we utilize the technical bilingual dictionary datasets provided by Maheshwari et al. (2024) which is a dataset curated from CSTT  [^3] https://cstt.education.gov.in/en dictionaries. The dataset consists of word-level translations from English to 6 Indian languages across 3 domains, viz., administrative, biotechnology, and chemistry, and has 9094 terms in the training data and 1285 in the test data for all domains combined. We obtained Sanskrit-based inputs for all data instances by applying our approach of generating Sanskrit-based additional inputs. We use chrF++ Popović (2017) as the evaluation metric for all the experiments under this task.
@@ 3.1 Experiments on the SWS Task
For the SWS task, we use word-level accuracy as the evaluation metric. To compare against Sandhan et al. (2022), we also calculate sentence level perfect match (PM) for SIGHUM and hackathon datasets. We utilize the pre-trained checkpoint of the base variant of the ByT5 model available via Huggingface  [^4] https://huggingface.co/google/byt5-base and fine-tune it over the UoH+SandhiKosh, SIGHUM dataset, and hackathon datasets as three separate experiments. For details about choice of ByT5 and Experiments with SWS Task (see Appendix B.2).
@@
3.2 Experiments on the Technical Term Translation Task
For this task, we have two experimental settings, both formulated as text-to-text translation. In the first setting, we train and test the NMT model NLLB Costa-jussà et al. (2022) over all 6 language pairs across 3 domains. In the second setting, we train the model on Hindi, Gujarati, and Tamil across 3 domains and test it over Marathi, Kannada, and Odia across the same domains, which can be considered as a zero-shot setting. In the baseline configuration for this task, the model is fed with English input only. In the configuration corresponding to the proposed method, the English input is augmented with additional Sanskrit-based input prepared as discussed in Section 2.2. We utilize the pre-trained 1.3B parameter checkpoint of the NLLB model available via Huggingface [^5] https://huggingface.co/facebook/nllb-200-1.3B and fine-tune it over the technical domain dictionary data for both experimental settings.
@@ Results.
Table 1 reports the comparison of chrF++ scores obtained by finetuning the NMT model with English-only input (NLLB) and with augmented input (NLLB+Sanskrit) under the first experimental setting. In Section 3.2 we analyze the performance of the model with and without additional input in a zero-shot setting. Across experiments, there’s a consistent performance gain with the lexically informed input. Our method archives an average improvement of 8.46 chrF++ scores. We also provide a detailed post-hoc analysis of the predictions in Section 3.3
Zero-Shot Translation
@@ 3.3 Post-hoc analysis
We followed the Human Post-hoc evaluation as per Maheshwari et al. (2024) for the same two additional languages as presented by them viz., Punjabi and Malayalam, using the same subset of input data and metrics. Our goal is to understand the practical utility of the generated lexicon in the respective languages and the extent to which they may be helpful in translating technical books from English to low-resource Indian languages. We achieved an R@1 score of 0.53 and 0.46 for Punjabi and Malayalam, respectively, compared to 0.51 and 0.38 scores obtained by LexGen. The R@3 score for Malayalam is 0.72, comparable to 0.71 for LexGen, while the score for Punjabi was slightly lower, at 0.92, compared to 0.95 for LexGen. We also present detailed analysis of the translation results by a comparative study of the outputs in both the input settings, i.e., with and without the Sanskrit-based augmented output (See Appendix A.3).
4 Conclusion
Limitations
Not all languages exhibit significant vocabulary overlap with Sanskrit, and in such cases, our proposed method may have limited applicability for lexicon generation.
Acknowledgments
We acknowledge IIT Bombay and BharatGen for providing resources and support for the project. Author Karthika acknowledges the PhD fellowship grant from the TCS Research Foundation. We also extend our appreciation to the reviewers for their valuable feedback.
Appendix A Appendix
A.1 Normalisation
A.2 sandhi
A.3 Post-hoc Analysis of Generated Technical Lexicons
@@
Appendix B Sanskrit Word Segmentation Task
B.1 Data - SWS Task
For the SWS task, following Dave et al. (2021) and Sandhan et al. (2022), we use three publicly available benchmark datasets, UoH corpus [^7] https://sanskrit.uohyd.ac.in/Corpus/ combined with the SandhiKosh dataset Bhardwaj et al. (2018), SIGHUM dataset Krishna et al. (2017), and hackathon dataset Krishnan et al. (2020). These datasets are carefully curated subsets of a larger corpus DCS Hellwig (2010). The UoH corpus+SandhiKosh dataset has 62273 and 15569 instances as train and test sets. For this dataset, we apply the pruning technique mentioned in Dave et al. (2021) to filter out invalid instances. The size of the training, validation, and test sets for the SIGHUM dataset are 97000, 3000, and 4200, respectively, and for the hackathon dataset, it is 90000, 10332, and 9963, respectively. Contemporary deep-learning methodologies have demonstrated enhanced performance when utilizing the SLP1 script for Sanskrit. Consequently, we have prepared all datasets in the SLP1 script to leverage these performance improvements.
Model LPA SPA
JNU - 8.1
UoH - 47.2
INRIA - 59.9
$D$-RNN 95.0 79.5
Sandhi Prakarana 92.3 86.8
ByT5 97.2 93.5
B.2 Experiment Details
Baselines.
For the experiments performed over the UoH+SandhiKosh dataset, we compare our method against Sandhi Prakarana Dave et al. (2021), DD-RNN Aralikatte et al. (2018), and 3 sandhi spitter tools viz (i) JNU Splitter Sachin (2007), (ii) UoH Splitter Kumar et al. (2010), and (iii) INRIA Sanskrit Heritage Reader Huet (2003); Goyal and Huet (2013). We reproduce and report the scores reported by Dave et al. (2021). For DD-RNN and the 3 sandhi tools, we report the scores reported in Aralikatte et al. (2018) and Dave et al. (2021). For the experiments performed over the SIGHUM and hackathon datasets, we compare our method against TransLIST Sandhan et al. (2022) and rcNN-SS Hellwig and Nehrdich (2018).
Results.
Tables 4 and 5 report the performance of our methodology compared with the baselines over the respective datasets. Table 4 shows that our methodology outperforms all other baselines in terms of both Location Prediction Accuracy (LPA) and Split Prediction Accuracy (SPA) with absolute gains of 4.86 and 6.72, respectively, on the UoH+SandhiKosh dataset. TransLIST Sandhan et al. (2022) utilizes a set of potential split candidates from SHR (referred to as LIST in their paper), which provides additional linguistic information for segmentation. Our model is not linguistically informed like this as we feed only the compound word to the model. Hence, our method is not strictly comparable with the results shown in row 2 of Table 5. Nevertheless, our method outperforms all other models on three out of four evaluation metrics when tested on hackathon dataset. On SIGHUM dataset, our method achieves competitive scores. Sandhan et al. (2022) also reported the performance of their model without the LIST module, as shown in row 3 (TransLIST). The model without the LIST step is more comparable to our setting and we outperform this result as well, while failing to outperform the scores in row 2. As a separate experiment, we provide SHR input to our model for SIGHUM data which outperforms TransLIST on PM metric achieving a PM score of 94.31.
Model SIGHUM Hackathon
P R F PM P R F PM
rcNN-$S$ 96.86 96.83 96.84 87.08 96.40 95.15 95.77 77.62
TransLIST 98.80 98.93 98.86 93.97 97.78 97.44 97.61 85.47
TransLIST - - - 86.10 - - - -
ByT5 98.68 98.42 98.53 93.78 97.58 97.71 97.63 87.7